An electron is placed midway between two parallel horizontal conducting plates 2.0 cm apart in air. What must be

the potential difference between the plates for the electrons to remain stationary? What

must be the sign of the charges in each plate?


1
Expert's answer
2022-03-30T13:49:33-0400

The equilibrium condition

"F_g=F_e\\
mg=qE\\
mg=qV/d"

Hence, the potential difference

"V=\frac{mgd}{q}\\
=\frac{9.1*10^{-31}*9.8*0.02}{1.6*10^{-19}}=1.11*10^{-12}\:\rm V"

The top plate should be positive.